OpenStudy (anonymous):

the solutions to the factored equation (x-2)(3x+4)=0 are?...is it -2 and 4?

jimthompson5910 (jim_thompson5910):

if (x-2)(3x+4)=0, then x-2 = 0 or 3x+4 = 0 solve each for x